The activation energy for the gas phase decomposition of ethyl acetate is 200 kJ.

CH3COOC2H5 --> CH3COOH + C2H4
The rate constant at 670 K is 8.41×10-4 /s. The rate constant will be 1.14×10-2 /s at _____ K.
Answer #1

ln(k2- ln k1) = Ea/R(1/T1 - 1/T2)

T1 = 670 K
k1 = 8.41 x 10^-4 s^-1

T2 = ? K
k2 = 1.14 x 10^-2 s^-1

R = 8.3145 J/K.mol (Gas constant)
Ea= 200 KJ OR 2.00 x 10^5 J/mol


ln 1.14 x 10^-2 - ln 8.41 x 10^-4 s^-1 = 2.00 x 10^5 J/mol / 8.3145 (1/670- 1/T2)

-4.47-(-7.08) = 24054.4 (1.5 x 10^-3 -1/T2 )

2.61= 24054.4 (1.5 x 10^-3 -1/T2 )

1.09 x 10^-4 =1.5 x 10^-3 -1/T2

1/T2 =1.5 x 10^-3 -1.09 x 10^-4

1/T2= 1.391 x 10^-3

T2= 1/1.391 x 10^-3

= 720 K